OpenStudy (derpisthewordorisit):

A box contains 12 green balls, 11 red balls, 7 white balls, and 10 black balls. What is the probability of drawing a black ball from this box? A. 1/10 B. 1/40 C. 1/4 D. 1/3

OpenStudy (derpisthewordorisit):

@Simple10

OpenStudy (kash_thesmartguy):

A. Confirm it.

OpenStudy (derpisthewordorisit):

No picture for this one.

OpenStudy (derpisthewordorisit):

Can you explain it, please?

OpenStudy (kash_thesmartguy):

It's C. Actually. Typo,

OpenStudy (derpisthewordorisit):

@Simple10 Is he right?

OpenStudy (anonymous):

Here is the work for this: Add all the balls. 12 + 11 + 7 + 10 = 40 So 10/40 is black which is eventually simplified down to 1/4. Your answer is C!

OpenStudy (derpisthewordorisit):

Thanks! can you help me with one more?

OpenStudy (kash_thesmartguy):

\[\frac{ 10 blackballs }{ 40 total balls }=\frac{ 1 }{ 4}\]

OpenStudy (anonymous):

Keep posting them and summoning me :| I love the medals ;)